# Not All Edges are Equally Robust: Evaluating the Robustness of Ranking-Based Federated Learning

### v1

@@ A.1 Algorithms
Algorithm 2 Edge-popup (EP) Algorithm
1:Number of local epochs $E$, training data $D$, initial weights $\theta^{w}$ and scores $\theta^{s}$, subnetwork size $k\%$, learning rate $\eta_{\text{ep}}$
2:for $e\in[1,E]$ do
3:     $B\leftarrow$ Split $D$ into $B$ batches
4:     for batch $b\in[B]$ do
5:         EP_Forward($\theta^{w},\theta^{s},k,b$)
6:         $\theta^{s}\leftarrow\theta^{s}-\eta_{\text{ep}}\nabla\ell[\theta^{s};b]$
7:     end for
8:end for
9:return $\theta^{s}$
10:function EP_Forward($\theta^{w},\theta^{s},k,b$)
11:     $m\leftarrow\text{sort}[\theta^{s}]$
12:     $t\leftarrow\text{int}[[1-k]\times\text{len}[m]]$
13:     $\theta_{p}=\theta^{w}\odot m$, where $m[:t]=0$; $m[t:]=1$
14:     return $\theta_{p}[b]$
15:end function
@@
11:         $\theta^{s}\leftarrow\text{Edge-PopUp}(E,D^{tr}_{u},\theta^{w},\theta^{s},k,\eta_{\text{ep}})$triangleright Client $u$ uses Algorithm 2 to train a supermask on its local training dataset
@@ A.2 Proof
Notation Description
$\theta^{w}$ model weight
$\theta^{s}$ score
$R_{u}$ Ranking for user $u$
$R^{2d}_{u}$ Expaned 2d ranking for user $u$
$\overline{\mathbb{E}}^{\text{in}}$ Top kedges selected by $U-m$ benign clients
$\overline{\mathbb{E}}^{\text{out}}$ Remaining edges
$\overline{\mathbb{W}}^{\text{in}}$ Reputations of top kedges selected by benign clients
$\overline{\mathbb{W}}^{\text{out}}$ Reputations of remaining edges
$\overline{w}_{\text{min}}$ Minimum reputation in $\overline{\mathbb{W}}^{\text{in}}$
$\overline{w}_{\text{max}}$ Maximum reputation in $\overline{\mathbb{W}}^{\text{out}}$
$\overline{w}_{j}$ Reputation for edge $e_{j}$ agreed by benign clients
$\widetilde{\mathbb{E}}^{\text{in}}$ Top kof edges selected by all clients after the attack
$\widetilde{\mathbb{E}}^{\text{out}}$ Remaining edges after the attack
$\widetilde{w}_{j}$ Reputation for edge $e_{j}$ agreed by all clients
TABLE V: Notation and description.
In this section, we detail the proof of Theorem 1. We will prove the upper bound of $\overline{w}_{v}$, and the method of proving the lower bound will be similar. We summarize the notations in Table V.
@@
TABLE VI: Model architectures
Architecture Layer Name Number of parameters
Conv2 Conv(64) 576
Conv(128) 73728
FC(256), FC(10) 1605632 or 2560
LeNet Conv(32) 288
Conv(64) 18496
FC(128), FC(10) 1605632 or 1280
Conv8 Conv(64), Conv(64) 38592
Conv(128), Conv(128) 221184
Conv(256), Conv(256) 884736
Conv(512), Conv(512) 3538944
FC(256), FC(256), FC(10) 592384
@@ Appendix B Additional Details on Experimental Setup
B.1 Datasets and Model Architectures
MNIST [41] consists of $70000$ grayscale images of handwritten digits, each with a size of $28\times 28$ pixels. The dataset is divided into $10$ classes representing the digits 0-9. For the model architecture, we experiment with Conv2 architecture given in Table VI. For the sparsity of the subnetwork, we noticed that selecting only 0.1 of the parameters could achieve the same accuracy as the original result, so we used sparsity as 0.1 for this task. For local training, we follow the configuration in [15]: each client trains for 2 epochs using an SGD optimizer with a learning rate of 0.4, momentum of 0.9, weight decay of 1e-4, and a batch size of 32.
CIFAR-10 [42] is a 10-class classification task consisting of $60000$ RGB images (50,000 for training and 10,000 for testing), each sized by $32\times 32$. For the model architecture, we experiment with Conv8 architecture given in Table VI. For the sparsity of the subnetwork, we followed FRL to select 0.5 of the model parameters as the subnetwork. For local training, we follow the configuration in [15]: each client trains for 5 epochs using an SGD optimizer with a learning rate of 0.4, momentum of 0.9, weight decay of 1e-4, and a batch size of 32.
Fashion-MNIST [43] is a 10-class fashion image for classification tasks, using a training set of $60000$ images and a testing set of $10000$ images with a resolution of $28\times 28$ pixels. We experiment with LeNet architecture given in Table VI. For local training, each client trains for 2 epochs using an SGD optimizer with a learning rate of 0.4, momentum of 0.9, weight decay of 1e-4, and a batch size of 32.
EMNIST [44] consists of $814255$ gray-scale images of numeric, upper, and lower case handwritten letters, with a total of 62 categories. We experiment with Conv2 architecture given in Table VI. For local training, each client trains for 5 epochs using an SGD optimizer with a learning rate of 0.4, momentum of 0.9, weight decay of 1e-4, and a batch size of 32.
CIFAR-100 [42] is a 100-class classification task consisting of $60000$ RGB images (50,000 for training and 10,000 for testing), each sized by $32\times 32$. For the model architecture, we experiment with Conv8 architecture given in Table VI. For the sparsity of the subnetwork, we followed FRL to select 0.5 of the model parameters as the subnetwork. For local training, we follow the configuration in [15]: each client trains for 5 epochs using an SGD optimizer with a learning rate of 0.4, momentum of 0.9, weight decay of 1e-4, and a batch size of 32.
B.2 Evaluated Attacks and Settings
@@
B.3 Evaluated Defenses and Settings
Krum [31] selects one local update that is close to its $U-m-2$ neighboring update in Euclidean distance as the global model; here, $m$ is an upper bound on the number of malicious clients and U is the number of clients selected for the current training.
Multi-Krum [31] is based on Krum to select $c$ updates such that $U-c>2m+2$, and then average the selected c updates as the global.
@@ B.3 Evaluated Defenses and Settings
FoolsGold [29] find clients with similar cosine similarity to be malicious, penalize their reputation, and return a weighted mean of the update weighed by their reputation.
@@ Appendix C Additional Experiment Results
Estimation Accuracy Fig. 14 shows the estimation accuracy during the FRL training in the non-IID setting, which still shows the desired result. The historical method achieves better performance than the alternative method.
Figure 14: Estimation accuracy during the FRL training in non-IID setting.
Attack Impact on Large Dataset Table VII shows the attack impact on the CIFAR10 dataset with VGG architecture and the CIFAR100 dataset with Conv8 architecture. From the table, we can see that our VEM achieves desirable attack impact.
FRL FABA FoolsGold Fang
CIFAR10 (VGG) 54.23 56.21 50.18 45.88
CIFAR100 (Conv8) 47.30 48.28 46.67 47.31
TABLE VII: Attack impact of VEM on the large dataset and model architecture.
Impact of Local Training Epoch Table VIII illustrates the attack impact of varying numbers of local training epochs, ranging from 2 to 10, on the MNIST dataset with FRL defense. We observe that as the number of local epochs increases, the attack impact slightly decreases. This is due to the reduction in prediction accuracy with more local epochs. However, the benign accuracy also declines, indicating that an increase in local training epochs negatively affects convergence ability. As shown in Fig 15, this phenomenon also appears in other datasets.
@@
TABLE VIII: Attack impact for the local training epoch conducted on MNIST dataset under FRL defense.
@@
Impact of Hyper-parameters
@@
More Defense Results. We also evaluate the False Positive Rate (FPR) and True Positive Rate (TPR) of the IBD defense under different values of hyper-parameter $\zeta$ in Eq. (7). As shown in Fig. 17, as the value of $\zeta$ increases, the FPR decreases while the TPR increases. However, it still has a relatively high FPR, e.g., more than 20% for the Fashion-MNIST dataset and around 20% for the CIFAR10 dataset, indicating that some benign updates are filtered out by the defense. Therefore, the IBD-integrated FRL is not yet sufficiently effective against our VEM attack.

### v2

@@ 6.2 Limitations and Future Work
Moreover, future research could explore certified robustness metrics. One possible evaluation approach is to determine the maximum number of malicious clients that the FRL system can tolerate under our attack. This requires further investigation into how changes in selected edges impact the model prediction results. Exploring this area could provide valuable insights for developing more robust FL systems.
@@ A.1 Algorithms
11:         $\theta^{s}\leftarrow\text{Edge-PopUp}(E,D^{tr}_{u},\theta^{w},\theta^{s},k,\eta_{\text{ep}})$triangleright Client $u$ uses Edge-PopUp to train a supermask on its local training dataset
@@ A.2 Proof
In this section, we detail the proof of Theorem 1. We will prove the upper bound of $\overline{w}_{v}$, and the method of proving the lower bound will be similar.
@@ Appendix B Additional Details on Experimental Setup
B.1 Detailed Settings
For Fig. 3, we conduct the experiment on MNIST (Conv2) and CIFAR-10 (Conv8) using the FRL framework with $k=50\%$ in a benign setting. In each round, we randomly select 25 clients for training and run the experiment for 1000 rounds. During each training round, we record the global ranking $R_{g}$ to calculate the reputation difference between each edge and the selection boundary. We then calculate the mean value of the reputation difference for each edge across 1000 rounds and use a bar graph (Fig. 3) to show its distribution. Additionally, we add a red dashed vertical line in Fig. 3 to indicate the ‘max adversary damage’. As the experiment is conducted in a benign setting, the ‘max adversary damage’ is calculated manually using $m(a_{n}-a_{1})$, where we assume $m=20\%$.
B.2 Datasets and Model Architectures
MNIST [41]: We Uses Conv2 architecture with 10% of parameters for sparsity. Each client trains for 2 epochs with SGD (LR=0.4, momentum=0.9, weight decay=1e-4, batch size=32).
CIFAR-10 [42]: We ues Conv8 architecture with 50% of parameters for sparsity. Each client trains for 5 epochs with SGD (same settings as MNIST).
Fashion-MNIST [43] We uses LeNet architecture with 50% of parameters for sparsity. Each client trains for 2 epochs with SGD (same settings as MNIST).
EMNIST [44]: We uses LeNet architecture with 50% of parameters for sparsity. Each client trains for 2 epochs with SGD (same settings as MNIST).
B.3 Evaluated Attacks and Settings
@@
B.4 Evaluated Defenses and Settings
Krum / Multi-Krum [31] selects one or more local updates that are close to its $U-m-2$ neighboring update in Euclidean distance as the global model; here, $m$ is an upper bound on the number of malicious clients and U is the number of clients selected for the current training.
@@ B.4 Evaluated Defenses and Settings
FoolsGold [29] finds clients with similar cosine similarity to be malicious, penalize their reputation, and return a weighted mean of the update weighed by their reputation.
@@ Appendix C Additional Experiment Results
Attack Impact on More Datasets. Table VI presents the impact of our attack on larger datasets, as well as three tabular datasets with 10% and 20% malicious rates under the FRL framework. From the table, we can see that our VEM achieves desirable attack impact in all cases.
Malicious CIFAR10 CIFAR100 Location30 Purchase100 Texas100
rate (VGG) (Conv8) (FC) (FC) (FC)
10% 44.21 40.54 88.01 95.65 51.56
20% 54.23 47.30 90.02 99.80 67.19
TABLE VI: Attack impact (%) of VEM on other datasets and model architectures with different malicious rates under FRL framework.
Impact of Local Training Epoch. Table VII illustrates the attack impact of varying numbers of local training epochs, ranging from 2 to 10, on the MNIST dataset with FRL defense. We observe that as the number of local epochs increases, the attack impact slightly decreases. This is due to the reduction in prediction accuracy with more local epochs. However, the benign accuracy also declines, indicating that an increase in local training epochs negatively affects convergence ability. As shown in Fig 15, this phenomenon also appears in other datasets.
@@
TABLE VII: Attack impact for the local training epoch conducted on MNIST dataset under FRL defense.
@@
Impact of Hyper-parameters.
@@
More Defense Discussion. Augmented defenses perform pre-processing steps before applying existing AGRs, enhancing the robustness of the FL system. We test our attack against the SOTA augmented defenses method: FoundationFL [47]. The experiment are conducted with 10% and 20% malicious rates, resulting in attack impacts of 66.25% and 72.50%, respectively. The reason FoundationFL fails to defend against our attack lies in its ‘synthetic updates generation’ process. This approach first selects updates that deviate significantly from extreme updates (i.e., those with large magnitudes) and uses them as synthetic updates. However, due to the nature of FRL, the server cannot accurately identify extreme updates. Consequently, the selection process becomes essentially random, resulting in the use of arbitrary updates as synthetic ones. This effect is functionally similar to increasing the number of selected clients per round. As shown in Fig. 11 in our manuscript, our attack remains unaffected by the client participation rate, enabling it to successfully bypass this defense.
Certified defenses provide provable security guarantees against poisoning attacks, making them more secure against advanced adversarial strategies. To evaluate the impact of our attack, we test it under FLCert [48]. We evaluate our attack against the FLCert defense on the CIFAR-10 dataset with 10% and 20% malicious rates, resulting in attack impacts of 83.33% and 86.53%, respectively. This outcome can be attributed to two key factors. First, in the early stages of model training, the model is not yet fully trained, leading to poor performance across all groups on the test samples. As training progresses, our attack gradually manipulates the global model, which subsequently influences each local model. As a result, the local models deviate from their optimal training trajectories, ultimately struggling on the test dataset.
Appendix D Meta-Review
D.1 Summary
The paper evaluates the robustness of Federated Ranking Learning (FRL), a framework for federated learning that uses rankings instead of gradients to improve communication efficiency and limit adversarial manipulation. While FRL mitigates some traditional vulnerabilities in federated learning, the study identifies specific edges in the network that are susceptible to targeted poisoning attacks. It introduces the Vulnerable Edge Manipulation (VEM) attack, which exploits these weaknesses using optimization techniques to maximize impact. Experimental results on benchmark datasets demonstrate that VEM significantly outperforms existing poisoning methods, with an attack impact up to 53.23%. Despite introducing a defense mechanism tailored to FRL, the proposed attack remains highly effective, highlighting critical vulnerabilities and emphasizing the need for robust and scalable federated learning frameworks.
D.2 Scientific Contributions
  • Identifies an Impactful Vulnerability
  • Provides a Valuable Step Forward in an Established Field
D.3 Reasons for Acceptance
  1. 1.
    First systemic work on studying the robustness of FRL, advancing the research in this space.
  2. 2.
    Reasonable threat model analyses that establish the practicality of the proposed attack.
  3. 3.
    Evaluations seem reasonable and outperform baselines by a comfortable margin.
D.4 Noteworthy Concerns
  1. 1.
    Certified robustness metrics remain un-evaluated, which is relevant to explore.